Determine whether each of the following functions from Z to Z is one to one and onto.


a.f(n)=n-1


b.f(n)=[n/2]


C.f(n)=n²+1

1
Expert's answer
2022-05-24T23:18:40-0400

a)


"f(n)=n-1, n\in \Z"

Let "f(n_1)=f(n_2), n_1, n_2\in \Z." It means that


"n_1-1=n_2-1"

"n_1=n_2"

The function "f(n)=n-1" is one-to-one from "\Z" to "\Z."

Let "y=n-1,y\in \Z." Then


"n=y+1"

We see that "\exist n\in \Z\ \forall y\in \Z."

The function "f(n)=n-1" is onto from "\Z" to "\Z."


The function "f(n)=n-1" is one-to-one and onto from "\Z" to "\Z."


(b)


"f(1)=\lceil1/2\rceil=1"

"f(2)=\lceil2/2\rceil=1"

"1\not=2, \ but\ f(1)=1=f(2)"

The function "f(n)=\lceil n/2\rceil" is not one-to-one from "\Z" to "\Z."


Let "y=k,y\in \Z." Take "n=2k, n\in \Z." Then


"\lceil n/2\rceil=\lceil 2k/2\rceil=\lceil k\rceil=k"

We see that "\forall y\in \Z\ \exist n\in\Z" such that "\lceil n/2\rceil=y."

The function "f(n)=\lceil n/2\rceil" is onto from "\Z" to "\Z."


The function "f(n)=\lceil n/2\rceil" is onto but is not one-to-one from "\Z" to "\Z."


(c)


"f(1)=(1)^2+1=2"

"f(-1)=(-1)^2+1=2"

"-1\not=1, \ but\ f(-1)=2=f(1)"

The function "f(n)=n^2+1" is not one-to-one from "\Z" to "\Z."


"\forall n\in \Z" the value of "f(n)=n^2+1" is non negative integer value.

Let "y=-1, y\in \Z."

We cannot find "n\in \Z" such that "f(n)=n^2+1=-1."

The function "f(n)=n^2+1" is not onto from "\Z" to "\Z."


The function "f(n)=n^2+1" is not ontoand is not one-to-one from "\Z" to "\Z."